OpenStudy (anonymous):

please help: New York states license plates contain 3 letters followed by 4 digits. The letters O and I are not allowed to be used on license plate. A) If letters and numbers are allowed to be replaced, how many different plates can be made?

OpenStudy (anonymous):

b) if the first letter has to be a vowel , how many less plates can be made? c: if the first place can be either a number or a letter, how many more license plates can be made than part a?

OpenStudy (kropot72):

a) The number of different permutations of n different things taken k at a time with repetitions is\[n ^{k}\] The number of permutations of the 24 allowable letters is \[24^{3}\] The number of permutations of the 10 digits 0 ~ 9 is \[10^{4}\] Each permutation of the letters can be taken with each of the possible permutations of digits. So the number of different plates that can be made is \[24^{3}\times 10^{4}=you\ can\ calculate\]

OpenStudy (anonymous):

so thats part a right?

OpenStudy (anonymous):

Okay i need help with part b and c?

OpenStudy (anonymous):

@kropot72 thanks for your help

OpenStudy (kropot72):

b) The vowels O and I are not allowed. If the letters start with the vowel A the number of permutations with the remaining 23 letters is 23^2. Similarly with the letters starting with the vowels E and U there are 23^2 permutations. So the total number of permutations of letters becomes \[3\times 23^{2}\] Each of these permutations of letters can be taken with each of the possible permutations of digits. So the total number of licence plates is \[3\times 23^{2}\times 10^{4}\] The reduction in the number of plates that can be made is \[(24^{3}\times 10^{4})-(3\times 23^{2}\times 10^{4})=you\ can\ calculate\]

OpenStudy (anonymous):

okay part c?

OpenStudy (kropot72):

c) Sorry, I can't understand this part of the question. Have you posted it exactly as the original? If not please post part c) again.

OpenStudy (kropot72):

Does part c) mean that the number plate would have an increase in the number of characters from 7 up to 8, with the first character being either a letter or a digit?

OpenStudy (anonymous):

if the first place can be either a number or a letter, how many more license plates can be made than in part a?

OpenStudy (anonymous):

hmmm... You want help on this last one?^^ Because it looks like your other questions got answered.

OpenStudy (anonymous):

yeah I got the others

OpenStudy (anonymous):

need help with the last ine!

OpenStudy (anonymous):

Okay! Just a second. :) Unless kropot here has it...

OpenStudy (anonymous):

@tafkas can u help me with another problem:) http://ms210q.schoolwires.net/cms/lib2/NY01001558/Centricity/Domain/150/HW%2098.pdf

OpenStudy (kropot72):

If there is an eighth character in the first position there are 10 digits that can be used and 24 allowable letters. In that case 10 + 24 = 34 times as many number plates can be made.

OpenStudy (anonymous):

34 times as many number plates can be made. thats the answer right?

OpenStudy (anonymous):

You need all three of these?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Okay, give me just a moment to be sure that *I* know how to do these, hee hee. :)

OpenStudy (anonymous):

okay! thanks!

OpenStudy (kropot72):

I confirm the answer to c) is 34 times as many number plates as in a).

OpenStudy (anonymous):

okay only #B okay thanks

OpenStudy (anonymous):

for the first one, you can pick any two ordered pairs and find the slope from there. Do you know the slope formula?

OpenStudy (anonymous):

yeah i didit that already and got slope is 2

OpenStudy (anonymous):

@kropot72 thanks!

OpenStudy (anonymous):

Now I need B please?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

Great! When they say, graph the data, they mean plot the points. They gave you points in part "A". Just plot them, and then you will see that the slope can be calculated visually, like so:

OpenStudy (anonymous):

hey! where'd my drawing go? >(

OpenStudy (anonymous):

so can I write on the grpah rise 4 run 2 thats how you can get slope

OpenStudy (anonymous):

do I need arrows?

OpenStudy (anonymous):

Great! Well, in theory, yes, because the arrows show that your line goes in opposite directions indefinitely (for infinity). The arrows wouldn't be there if you had a specific starting point, but you don't in this question.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

You're welcome! :D

OpenStudy (anonymous):

Show how to write an equation for the table and graph. Verify that the equation works with the table and the graph. is y=2x right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Well, yes! :D

OpenStudy (anonymous):

Great work! ;)